OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 16 less than or greater to 11 and x + 16 greater than or equal to 3 A. 11 greater than or equal to x + 16 greater than or equal to 3 B. -5 greater than or equal to x greater than or equal to -13 C. A number line with a closed circle on -13, a closed circle on -5, and shading in between. D. A number line with a closed circle on -13, shading to the left, and a closed circle on -5, shading to the right.

OpenStudy (amistre64):

it would help out alot if you wrote this in some kind of mathmatical notation

OpenStudy (amistre64):

<= means less than or equal to; >= means greater than or equal to

OpenStudy (anonymous):

D as the number range is between -13 and -5 and D is the opposite of this range i.e. \[-13 \ge x \le-5\]

OpenStudy (anonymous):

Thank You Kevsturge!

OpenStudy (anonymous):

D is the answer but I messed up the latex so \[-13 \le x \le -5\]